Carefully draw two isosceles triangles. Draw them so that each triangle has two sides of length 10 but one triangle has a third side of length 12 and the other triangle has a third side of length 16. Which of these two triangles has the largest area? Please show your work.
Answer by richwmiller(17219)

The triangle with base 12 has a height of 8.
and each half of the isosceles triangle is a right triangle of sides 10, 6 and 8
1/2 bh=1/2*12*8=1/2*4*3*2*4
The triangle with base 16 has a height of 6.
and each half of the isosceles triangle is a right triangle of sides 10, 6 and 8
1/2 bh=1/2*16*6=1/2*4*4*3*2
They are equal in area